# Toward axion signal extraction in semiconductor spin qubits via spectral engineering

### v2

@@ I Introduction
The first model for the axion field, established in 1977, provides a new symmetry mechanism that solves the strong CP problem in quantum chromodynamics [1, 2, 3]. However, the weak coupling between axions and other Standard Model particles makes the direct experimental detection of axion-related signals difficult. Moreover, the signal is often overwhelmed by environmental and measurement noise. Earlier detection proposals considered axion–electromagnetic field couplings, demonstrating that axion fields can be incorporated into Maxwell’s equations by adding extra source terms. This reformulation transforms elusive axion fields into detectable electromagnetic signals [4]. Building on this idea, various experimental and theoretical studies over the past decades have significantly improved the constraints on axion parameter boundaries, including the axion decay constant ($f_{a}$), the axion–photon coupling constant ($g_{a\gamma\gamma}$), and the axion–electron coupling constant ($g_{ae}$) [5, 6, 7]. A typical experimental architecture relies on a resonant microwave cavity within a strong superconducting magnet to detect axion–photon conversion at cryogenic temperatures [8, 9, 10, 11]. Pioneering works, such as the transmon-based Qub-IT approach, use a high-$Q$ superconducting cavity operating in GHz range, combined with quantum non-demolition single-photon counting via fixed-frequency transmon qubits operated below 20 mK [12]. While promising, this method requires a precise design of the cavity resonant frequency, which may limit the frequency bandwidth for axion searches.
@@
In this manuscript, we first introduce the axion model, followed by the Hamiltonian of the electron spin qubit platform. To relate the general axion framework to a concrete realization, we adopt the Dine–Fischler–Srednicki–Zhitnitsky (DFSZ) model as our specific axion candidate [36, 37]. The DFSZ axion model arises from a two-Higgs-doublet extension of the Standard Model, in which the Peccei–Quinn symmetry is carried by both quarks and leptons via Yukawa interactions. The tree-level axion–electron coupling is given by:
@@
Then, we can analyze the qubit–axion interactions in the non-relativistic limit ($\dot{a}^{2}\gg(\nabla a)^{2}$) and derive an effective qubit Hamiltonian. From this effective Hamiltonian, we extract the energy spectrum variations due to the magnetic field induced by the axion field, and use standard density-matrix (Lindblad) equations to compute the spin-polarization dynamics, which is the experimentally measurable quantity. For the postprocessing, this subtly modulated interaction signal is passed through a band-pass filter centered on the expected sideband frequencies to constrain the analysis bandwidth and reject out-of-band noise. While this filtering does not reduce the fundamental noise spectral density at the sidebands—and thus does not alter the intrinsic frequency-domain SNR—it confines noise power to a narrower band, making the two modulation sidebands more readily distinguishable against a suppressed out-of-band noise background.
@@ II.3 Noise simulation and signal processing protocol
To make the best use of the potential of quantum sensing in quantum dot array platform, we discuss a possible blueprint for the experimental setup to extend the lower boundary of the qubit-axion couplings and the effective magnetic field. A starting point is the dispersive coupling (with $g_{j,\text{bare}}\ll|\Delta_{j}|$, $\Delta_{j}=\omega_{01}^{(j)}-\omega_{c}$, and $\omega_{01}^{(j)}$ the Larmor frequency of the $j$-th qubit) between $N$ semiconductor spin qubits and a single mode of a high-$Q$ microwave resonator at frequency $\omega_{c}$ [50, 51]. Although current CMOS technology is still maturing toward large-scale entanglement and should not be over-interpreted, recent proposals have demonstrated industrial-scale quantum-dot arrays [25, 48].
The simplified multi-qubit–resonator Hamiltonian in the dispersive regime is
@@
where $a$ ($a^{\dagger}$) is the cavity-mode annihilation (creation) operator, and $\chi_{j}\equiv g_{j,\text{bare}}^{2}/\Delta_{j}$.
@@
whose time-dependent modulations encode the axion–qubit interaction. After the interaction period, the RF reflectometry signal from the gate-based LC resonator on the CMOS quantum dot array is routed into a cryogenic CMOS readout ASIC, which co-packages: 1. a two-stage SiGe/CMOS LNA operating at the 4K stage with a noise temperature $T_{N}\lesssim 4\,$K (with future goals of $T_{N}\lesssim 1\,$K). 2. a cryo-CMOS IQ down-converter (mixer + on-chip narrow-band IF filters) referenced to the qubit Larmor frequency $\omega_{01}$. 3. and an in-pixel Skipper-inspired ADC (65nm CMOS) that performs $M$ non-destructive charge-sampling (“pile-up") in the analogue domain before a 10-bit SAR conversion at about $100$ kS/s, achieving sub–single-electron noise ($0.48e^{-}$rms) at $T\approx 100$ K for spin–to–charge conversion. The serialized digital output of this ASIC is sent over a low-power link to a room-temperature DAQ system, where the further process is applied. one branch reconstructs and FFT-processes the demodulated time-domain waveform for broadband spectral analysis, and the other ingests the multi-sampled ADC stream for high-resolution, narrowband detection of spin-resonance shifts. By integrating the LNA, down-converter and multi-sampling ADC on a single cryo-CMOS chip directly adjacent to the quantum dot sensor (rather than placing the ADC at room temperature), parasitic loss and thermal noise are minimized while preserving sub–single-electron resolution.
@@ III Results and discussion
Although our present sensitivity (Fig. 3) does not yet fully cover the classical DFSZ band for QCD axions, the consideration of CMOS quantum-dot spin platform remains highly significant for several reasons. First, any ALPs with an electron coupling $g_{ae}$ above our sensitivity threshold can be probed, including models with suppressed photon couplings or mass scales outside the QCD axion window. Second, the cryo-CMOS readout ASIC, co-packaged with the quantum-dot array, can be tiled to $10^{4}$$10^{6}$ pixels on a single 300 mm wafer, enabling future detectors with $N\approx 10^{8}$$10^{10}$ spins. Here we also sketch a possible roadmap to further improve the performance. 1. Advances in materials and spin control are pushing quantum-dot $T_{2}$ from the microsecond to the millisecond scale. 2. New high-$Q$ three-dimensional and photonic-crystal resonator designs promise field-enhancement factors $Q\gtrsim 10^{6}$. 3. Next-generation amplifier like JPAs/TWPAs and Skipper-inspired CMOS ADCs are on track to achieve noise temperatures $T_{N}<0.1$ K and sub–single-electron resolution at about 100 kS/s. With these developments, we anticipate that a future device incorporating $N\gtrsim 10^{8}$ spins, millisecond coherence, and megahertz-bandwidth high-$Q$ readout will extend sensitivity well into—and ultimately beyond the DFSZ parameter region.
@@
A realistic noise model (cf. Table 2) is indispensable for our simulations because spin-qubit experiments inevitably experience a combination of broadband electronic noise, such as low-frequency $1/f$ fluctuations, quantization errors, discrete two-level system jumps, thermal drifts, and laboratory line-frequency pickup [13, 17]. Accordingly, we set the white noise amplitude to $1\times 10^{-3}$ and the pink noise amplitude to $2\times 10^{-2}$ (with $f_{\min}=1$ Hz) to replicate measured spectral densities of room-temperature amplifiers and device flicker noise; the readout noise parameters ($\sigma=2\times 10^{-3}$, $p_{\text{spike}}=5\times 10^{-3}$) to model digitization uncertainty and rare transient glitches in the data acquisition chain; the telegraph noise amplitude of $5\times 10^{-4}$ and drift amplitude of $2\times 10^{-4}$ to account for charge fluctuators and slow thermal variations in cryogenic components; and a modulated AC term ($A=1\times 10^{-4}$ at $f_{\text{mod}}=50$ Hz) to include residual coupling to the power-line environment. And the Johnson–Nyquist noise has the power density of $4k_{B}TR$ with $T=300$ K. These empirically motivated parameter choices ensure that our sensitivity estimates and SNR predictions accurately reflect the complex noise landscape encountered in high-precision axion search experiments.
@@
Fig. 6 presents the noisy simulation results for the spin polarization expectation value $\expectationvalue{x}$. After application of the filtering procedure (see Fig. 6(a)), a well-defined oscillatory component emerges at the theoretically predicted modulation frequency interval, exhibiting close quantitative agreement with the corresponding noiseless prediction. In Fig. 6(c), it demonstrates the PSD for noise only, which is higher at the low frequency but still non-zero and negligible at the desired signal frequency.
As shown in Fig. 6(d), the normalized cumulative power exhibits discrete, step-like increases not only at the qubit fundamental resonance frequency $f_{0}$ but also at the sideband frequencies $f_{0}\pm f_{\text{mod}}$ instead (indicated by the vertical dashed lines). These inflection points coincide precisely with the theoretically predicted positions of the modulation-induced sidebands corresponding to the peaks with vertical green lines in Fig. 6(b), demonstrating that a significant fraction of the desired signal energy is also carried by the axion-driven sideband components rather than being only confined solely to the qubit itself. Consequently, the modulation power appears predominantly in these sidebands rather than at the un-modulated resonance.
@@
Note that in z-axis signal processing, applying a causal band-pass filter to a signal containing abrupt transitions or broadband noise will convolve the data with the filter’s oscillatory impulse response, producing decaying sinusoidal ringing artifacts-even when the actual underlying signal is zero, since the $\expectationvalue{\sigma_{z}}$ should always be zero without the noise unlike the $\expectationvalue{\sigma_{x}}$. And in this case, the signal level is below the noise floor, which means there is no axion influence.

### v3

@@ I Introduction
The first model for the axion field, established in 1977, provides a new symmetry mechanism that solves the strong CP problem in quantum chromodynamics [1, 2, 3]. However, the weak coupling between axions and other Standard Model particles makes the direct experimental detection of axion-related signals difficult. Moreover, the signal is often overwhelmed by environmental and measurement noise. Earlier detection proposals considered axion–electromagnetic field couplings, demonstrating that axion fields can be incorporated into Maxwell’s equations by adding extra source terms. This reformulation transforms elusive axion fields into detectable electromagnetic signals [4]. Building on this idea, various experimental and theoretical studies over the past decades have significantly improved the constraints on axion parameter boundaries, including the axion decay constant ($f_{a}$), the axion–photon coupling constant ($g_{a\gamma\gamma}$), and the axion–electron coupling constant ($g_{ae}$) [5, 6, 7]. A typical experimental architecture relies on a resonant microwave cavity within a strong superconducting magnet to detect axion–photon conversion at cryogenic temperatures [8, 9, 10, 11]. Pioneering works, such as the transmon‐based Qub‐IT approach, use a high‐$Q$ superconducting cavity operating in GHz range, combined with quantum non-demolition single‐photon counting via fixed‐frequency transmon qubits operated below 20 mK [12]. While promising, this method requires a precise design of the cavity resonant frequency, which may limit the frequency bandwidth for axion searches.
@@
In this manuscript, we first introduce the axion model, followed by the Hamiltonian of the electron spin qubit platform. To relate the general axion framework to a concrete realization, we adopt the Dine–Fischler–Srednicki–Zhitnitsky (DFSZ) model as our specific axion candidate [36, 37]. The DFSZ axion model arises from a two-Higgs-doublet extension of the Standard Model, in which the Peccei–Quinn symmetry is carried by both quarks and leptons via Yukawa interactions. The tree‐level axion–electron coupling is given by:
@@
Then, we can analyze the qubit–axion interactions in the non-relativistic limit ($\dot{a}^{2}\gg(\nabla a)^{2}$) and derive an effective qubit Hamiltonian. From this effective Hamiltonian, we extract the energy spectrum variations due to the magnetic field induced by the axion field, and use standard density-matrix (Lindblad) equations to compute the spin-polarization dynamics, which is the experimentally measurable quantity. For the postprocessing, this subtly modulated interaction signal is passed through a band‐pass filter centered on the expected sideband frequencies to constrain the analysis bandwidth and reject out‐of‐band noise. While this filtering does not reduce the fundamental noise spectral density at the sidebands—and thus does not alter the intrinsic frequency‐domain SNR—it confines noise power to a narrower band, making the two modulation sidebands more readily distinguishable against a suppressed out‐of‐band noise background.
@@ II.3 Noise simulation and signal processing protocol
To make the best use of the potential of quantum sensing in quantum dot array platform, we discuss a possible blueprint for the experimental setup to extend the lower boundary of the qubit-axion couplings and the effective magnetic field. A starting point is the dispersive coupling (with $g_{j,\text{bare}}\ll|\Delta_{j}|$, $\Delta_{j}=\omega_{01}^{(j)}-\omega_{c}$, and $\omega_{01}^{(j)}$ the Larmor frequency of the $j$-th qubit) between $N$ semiconductor spin qubits and a single mode of a high-$Q$ microwave resonator at frequency $\omega_{c}$ [50, 51]. Although current CMOS technology is still maturing toward large-scale entanglement and should not be over-interpreted, recent proposals have demonstrated industrial‐scale quantum‐dot arrays [25, 48].
The simplified multi‐qubit–resonator Hamiltonian in the dispersive regime is
@@
where $a$ ($a^{\dagger}$) is the cavity‐mode annihilation (creation) operator, and $\chi_{j}\equiv g_{j,\text{bare}}^{2}/\Delta_{j}$.
@@
whose time-dependent modulations encode the axion–qubit interaction. After the interaction period, the RF reflectometry signal from the gate‐based LC resonator on the CMOS quantum dot array is routed into a cryogenic CMOS readout ASIC, which co‐packages: 1. a two‐stage SiGe/CMOS LNA operating at the 4K stage with a noise temperature $T_{N}\lesssim 4\,$K (with future goals of $T_{N}\lesssim 1\,$K). 2. a cryo‐CMOS IQ down‐converter (mixer + on‐chip narrow-band IF filters) referenced to the qubit Larmor frequency $\omega_{01}$. 3. and an in‐pixel Skipper‐inspired ADC (65nm CMOS) that performs $M$ non‐destructive charge‐sampling (“pile‐up") in the analogue domain before a 10‐bit SAR conversion at about $100$ kS/s, achieving sub–single‐electron noise ($0.48e^{-}$rms) at $T\approx 100$ K for spin–to–charge conversion. The serialized digital output of this ASIC is sent over a low‐power link to a room‐temperature DAQ system, where the further process is applied. one branch reconstructs and FFT‐processes the demodulated time‐domain waveform for broadband spectral analysis, and the other ingests the multi‐sampled ADC stream for high‐resolution, narrowband detection of spin‐resonance shifts. By integrating the LNA, down‐converter and multi‐sampling ADC on a single cryo‐CMOS chip directly adjacent to the quantum dot sensor (rather than placing the ADC at room temperature), parasitic loss and thermal noise are minimized while preserving sub–single‐electron resolution.
@@ III Results and discussion
Although our present sensitivity (Fig. 3) does not yet fully cover the classical DFSZ band for QCD axions, the consideration of CMOS quantum‐dot spin platform remains highly significant for several reasons. First, any ALPs with an electron coupling $g_{ae}$ above our sensitivity threshold can be probed, including models with suppressed photon couplings or mass scales outside the QCD axion window. Second, the cryo-CMOS readout ASIC, co-packaged with the quantum-dot array, can be tiled to $10^{4}$$10^{6}$ pixels on a single 300 mm wafer, enabling future detectors with $N\approx 10^{8}$$10^{10}$ spins. Here we also sketch a possible roadmap to further improve the performance. 1. Advances in materials and spin control are pushing quantum-dot $T_{2}$ from the microsecond to the millisecond scale. 2. New high-$Q$ three-dimensional and photonic-crystal resonator designs promise field-enhancement factors $Q\gtrsim 10^{6}$. 3. Next-generation amplifier like JPAs/TWPAs and Skipper-inspired CMOS ADCs are on track to achieve noise temperatures $T_{N}<0.1$ K and sub–single-electron resolution at about 100 kS/s. With these developments, we anticipate that a future device incorporating $N\gtrsim 10^{8}$ spins, millisecond coherence, and megahertz-bandwidth high-$Q$ readout will extend sensitivity well into—and ultimately beyond the DFSZ parameter region.
@@
A realistic noise model (cf. Table 2) is indispensable for our simulations because spin-qubit experiments inevitably experience a combination of broadband electronic noise, such as low‐frequency $1/f$ fluctuations, quantization errors, discrete two‐level system jumps, thermal drifts, and laboratory line‐frequency pickup [13, 17]. Accordingly, we set the white noise amplitude to $1\times 10^{-3}$ and the pink noise amplitude to $2\times 10^{-2}$ (with $f_{\min}=1$ Hz) to replicate measured spectral densities of room‐temperature amplifiers and device flicker noise; the readout noise parameters ($\sigma=2\times 10^{-3}$, $p_{\text{spike}}=5\times 10^{-3}$) to model digitization uncertainty and rare transient glitches in the data acquisition chain; the telegraph noise amplitude of $5\times 10^{-4}$ and drift amplitude of $2\times 10^{-4}$ to account for charge fluctuators and slow thermal variations in cryogenic components; and a modulated AC term ($A=1\times 10^{-4}$ at $f_{\text{mod}}=50$ Hz) to include residual coupling to the power‐line environment. And the Johnson–Nyquist noise has the power density of $4k_{B}TR$ with $T=300$ K. These empirically motivated parameter choices ensure that our sensitivity estimates and SNR predictions accurately reflect the complex noise landscape encountered in high‐precision axion search experiments.
@@
Fig. 6 presents the noisy simulation results for the spin polarization expectation value $\expectationvalue{x}$. After application of the filtering procedure (see Fig. 6(a)), a well‐defined oscillatory component emerges at the theoretically predicted modulation frequency interval, exhibiting close quantitative agreement with the corresponding noiseless prediction. In Fig. 6(c), it demonstrates the PSD for noise only, which is higher at the low frequency but still non-zero and negligible at the desired signal frequency.
As shown in Fig. 6(d), the normalized cumulative power exhibits discrete, step‐like increases not only at the qubit fundamental resonance frequency $f_{0}$ but also at the sideband frequencies $f_{0}\pm f_{\text{mod}}$ instead (indicated by the vertical dashed lines). These inflection points coincide precisely with the theoretically predicted positions of the modulation‐induced sidebands corresponding to the peaks with vertical green lines in Fig. 6(b), demonstrating that a significant fraction of the desired signal energy is also carried by the axion‐driven sideband components rather than being only confined solely to the qubit itself. Consequently, the modulation power appears predominantly in these sidebands rather than at the un-modulated resonance.
@@
Note that in z-axis signal processing, applying a causal band‐pass filter to a signal containing abrupt transitions or broadband noise will convolve the data with the filter’s oscillatory impulse response, producing decaying sinusoidal ringing artifacts-even when the actual underlying signal is zero, since the $\expectationvalue{\sigma_{z}}$ should always be zero without the noise unlike the $\expectationvalue{\sigma_{x}}$. And in this case, the signal level is below the noise floor, which means there is no axion influence.